# CART: Caltech Aerial RGB-Thermal Dataset in the Wild

### v1

@@ Abstract
We present the first publicly available RGB-thermal dataset designed for aerial robotics operating in natural environments. Our data-set captures a variety of terrains across the continental United States, including rivers, lakes, coastlines, deserts, and forests, and consists of synchronized RGB, long-wave thermal, global positioning, and inertial data. Furthermore, we provide semantic segmentation annotations for 10 classes commonly encountered in natural settings in order to facilitate the development of perception algorithms robust to adverse weather and nighttime conditions. Using this dataset, we propose new and challenging benchmarks for thermal and RGB-thermal semantic segmentation, RGB-to-thermal image translation, and visual-inertial odometry. We present extensive results using state-of-the-art methods and highlight the challenges posed by temporal and geographical domain shifts in our data. Dataset and accompanying code will be provided at https://github.com/aerorobotics/caltech-aerial-rgbt-dataset.
@@
**footnotetext: These authors contributed equally to this work
Figure 1: Left: Our dataset is the first dataset designed to improve thermal semantic perception for field robotics. Right: Our dataset provides new thermal and RGB-thermal benchmarks for field robotics and computer vision algorithms, encompassing (a) semantic segmentation, (b) image translation, and (c) motion tracking.
@@ 1 Introduction
Current field robots rely predominantly on sensors such as visual cameras, lidar, and radar to perceive their surroundings [16, 18]. While these sensors enhance robustness of downstream vision algorithms, their performance degrades in low-light and adverse weather (snow, fog, rain) conditions [68]. In contrast, thermal cameras exploit long-wave infrared wavelengths to capture emitted heat, offering dense, visual information even in conditions in which other methods struggle [19]. Recently, thermal cameras have been used in multimodal perception algorithms for autonomous vehicle applications and are increasingly being explored for field robotic applications to enable nighttime autonomy [36, 74, 69, 11, 76, 48, 12]. However, successful integration of thermal imagery requires extensive datasets across diverse settings, preventing widespread adoption in field robotics.
@@
Existing thermal datasets primarily focus on urban environments for autonomous driving applications (Fig. 1). They typically comprise of thermal-only imagery [74, 36] or RGB-Thermal (RGB-T) pairs [17, 22, 69, 40], with some including global positioning (GPS/GNSS) and inertial measurements (IMU) for visual-inertial odometry (VIO) and simultaneous localization and mapping (SLAM) [76, 11, 60]. While comprehensive, these datasets rarely extend beyond urban areas. They lack data depicting natural settings like rivers and forests, which are typical operating areas for field robotics that perform coastline mapping and bathymetry [4] or monitor activity during forest fires [30].
@@
Due to the lack of thermal benchmarks for algorithms like semantic segmentation and SLAM, field robots cannot easily operate at night or in adverse conditions, requiring online learning [70, 34] or unsupervised domain adaptation to compensate [20, 32, 69] which still require labeled data for evaluation. As such, enabling thermal perception for field robotics requires collecting and annotating field-specific data from scratch. However, curating such a dataset is more complex than doing so for RGB: First, thermal data is hard to crowdsource and web-scrape due to the cost and nicheness of thermal cameras. Hence, acquiring a dataset for algorithm development requires physically going to distinct field environments, increasing time and cost. Second, operating most field robotics, especially uninhabited aerial vehicles (UAV), usually require special permits unique to each location-of-capture. This further complicates the creation of a comprehensive thermal dataset for field robotics research.
@@
In this work, we present the first dataset targeting thermal scene perception for field robotics and for wider use by the computer vision community. The dataset consists of oblique-facing imagery captured from a UAV, supplemented with image sets captured at ground level, and focuses on littoral settings within various desert, forest, and coastal environments across the United States. We contribute new benchmarks for thermal and RGB-T semantic segmentation, RGB-T image translation, and VIO/SLAM algorithms, with unique challenges characterized by temporal and geographical domain shift for learning-based methods and periodic feature sparsity for motion tracking algorithms.
@@ 2.1.1 Thermal/RGB-T benchmarks for urban robotics
Current urban thermal and RGB-T datasets primarily focus on autonomous vehicle (AV) technology and surveillance applications. AV-related datasets cover tasks such as object detection [17, 40, 11], semantic segmentation [36, 22, 69, 74], and localization [76, 11, 60]. Some include data from other spectra, benchmarking not only thermal algorithms but also RGB-T [40, 11, 22, 69] and other multispectral models. In contrast, datasets for surveillance applications use fixed cameras [29] and UAVs [65] to detect personnel, vehicles, and other objects. Despite many urban thermal benchmarks, they cannot be directly used to develop algorithms for robots operating in natural environments due to the urban/non-urban domain gap.
@@ 2.1.2 Thermal/RGB-T benchmarks for field robotics
Datasets for robots operating in non-urban environments remain limited. MassMIND [48] and PST900 [61] benchmark thermal semantic segmentation for maritime and subterranean robotic environments, respectively. WIT-UAS [30] provides UAV-borne thermal images for object detection in wildfire-prone environments. [34] proposes an aerial RGB-T dataset for thermal water segmentation in littoral areas but only provides 1272 labels consisting of a single water class. Benchmarks like [3, 7] depict natural environments but are targeted towards wildlife tracking. To address this gap, we present an aerial RGB-T dataset which can be used to benchmark multiple tasks, including semantic segmentation, RGB-T image translation, and localization, across diverse natural environment robotic applications. Specifically, we build upon the dataset from [34] by extending the number of semantic classes from 1 to 10 and adding additional RGB-T data captured from air and ground, resulting in a total of 4195 semantic segmentation annotations.
@@ 2.1.3 UAV-based semantic segmentation datasets
Aside from [34], existing UAV-captured semantic segmentation datasets [47, 10, 44, 6, 46] cover urban environments. Wth the exception of [46] which provides 400 thermal image samples, all of these datasets contain only RGB data. Although they provide images with similar view angles as in our dataset, the RGB-T modality gap and their focus on urban semantic content makes them impractical for field robotic applications.
@@ 2.2.4 Thermal VIO and SLAM
3 The Aerial RGB-Thermal Dataset
@@ 3.3.3 Geographical split
    Region-based: This split partitions data based on the general area of the United States that the data was captured. The regions include California (CA), Kentucky (KY), and North Carolina (NC), with train/val/test splits created in the same manner as the terrain split.
@@ 4 Experiments
Model
Bare ground
Rocky terrain
Develop. struct.
Road
Shrubs
Trees
Sky
Water
Vehicles
Person
Stuff
Object
All
Orin (GPU)
NUC (CPU)
FLOPS (G)
FastSCNN [52]
0.807
0.907
0.728
0.631
0.694
0.776
0.956
0.965
0.384
0.058
0.808
0.221
0.690
124.2 85.6 1.0
MobileNetV3-S 0.75 [25]
0.802
0.914
0.676
0.633
0.704
0.814
0.955
0.976
0.383
0.111
0.809
0.247
0.697
75.2
24.0
3.3
MobileNetV3-L 0.75
0.801
0.914
0.701
0.615
0.713
0.793
0.959
0.976
0.432
0.148
0.809
0.290
0.705
59.0
15.5
4.8
MobileViTV2 0.50 [45]
0.822
0.913
0.759
0.650
0.694
0.781
0.954
0.970
0.222
0.030
0.818
0.126
0.680
60.1
10.7
5.5
EfficientViT-B0 [5]
0.825
0.917
0.777
0.618
0.710
0.793
0.961
0.965
0.495
0.191
0.821
0.343 0.725
51.2
5.5
3.9
EfficientNet-Lite0 [66]
0.819
0.908
0.769
0.601
0.709
0.805
0.962
0.968
0.380
0.079
0.818
0.229
0.700
47.7
12.8
7.2
EfficientNet-Lite2
0.825
0.920
0.765
0.614
0.698
0.805
0.958
0.971
0.386
0.118
0.820
0.252
0.706
38.9
11.2
9.4
Segformer-B0 [72]
0.804
0.910
0.690
0.624
0.696
0.788
0.960
0.970
0.195
0.000
0.805
0.097
0.664
38.3
5.9
10.2
Segformer-B1
0.814
0.909
0.773
0.603
0.713
0.799
0.960
0.965
0.366
0.000
0.817
0.183
0.690
30.3
3.5
19.9
ResNet18 [23]
0.810
0.905
0.766
0.618
0.711
0.807
0.959
0.966
0.431
0.158
0.818
0.294
0.713
69.3
12.4
22.4
ResNet50
0.819
0.916
0.747
0.658
0.711
0.799
0.961
0.974
0.361
0.182
0.823
0.272
0.713
29.9
6.1
45.4
ResNeXt50 [73]
0.825
0.919
0.773
0.653
0.709
0.794
0.964
0.976
0.436
0.137
0.827
0.287
0.719
23.0
5.6
45.5
ConvNext-T [43]
0.799
0.909
0.719
0.608
0.706
0.808
0.961
0.970
0.363
0.003
0.810
0.183
0.685
25.8
3.3
47.4
ConvNext-S
0.810
0.913
0.771
0.603
0.718
0.810
0.965
0.964
0.492
0.048
0.819
0.270
0.709
19.1
2.3
75.0
ConvNext-B
0.810
0.921
0.697
0.611
0.706
0.803
0.963
0.972
0.348
0.112
0.811
0.230
0.694
14.6
1.4
130.5
ConvNext-B (CLIP) [54]
0.813
0.918
0.683
0.632
0.718
0.813
0.965
0.972
0.517
0.137
0.814
0.327
0.717
14.6
1.4
130.4
ConvNext-B (CLIP) ❄
0.773
0.887
0.713
0.487
0.656
0.756
0.946
0.951
0.158
0.019
0.771
0.089
0.635
14.6
1.4
130.4
DINOv2 [49] (linear head)
0.800
0.907
0.681
0.606
0.693
0.796
0.956
0.967
0.375
0.079
0.801
0.227
0.686
15.2
2.1
DINOv2 (linear head) ❄
0.705
0.844
0.619
0.419
0.558
0.725
0.922
0.920
0.356
0.158
0.714
0.257
0.623
15.2
2.1
DINOv2 (nonlin. head)
0.810
0.916
0.708
0.635
0.700
0.819
0.959
0.973
0.399
0.119
0.815
0.259
0.704
15.2
2.1
DINOv2 (nonlin. head) ❄
0.796
0.903
0.732
0.606
0.691
0.794
0.954
0.963
0.471
0.149
0.805
0.310
0.706
15.2
2.1
${}^{\dagger}$ [50]
0.755
0.867
0.635
0.576
0.643
0.653
0.787
0.947
0.234
0.024
0.733
0.129
0.613
11.1
1.4
100.5
${}^{\dagger}$hermal-specific model  ❄ Frozen encoder
@@ 4.1.3 Geographically-partitioned benchmarks
(Tab. 1) on geographically-partitioned sets to evaluate intra-thermal generalization. Given the cost and effort of collecting diverse aerial thermal data of field settings, we aim to determine the least amount of annotated data required to effectively deploy. We performed two experiments using EfficientViT-B0: one on generalization across geographic regions and another across differing terrain types.
@@
River
Lake
Coast
River
0.668
0.196
0.117
0.327
Lake
0.196
0.364
0.061
0.207
Coast
0.127
0.119
0.522
0.256
All
0.646
0.415
0.493
0.518
All + ground
0.658
0.416
0.520
0.532
@@ 4.1.4 Temporally-partitioned benchmark
Model
Bare ground
Rocky terrain
Devel. struct.
Road
Shrubs
Trees
Sky
Water
Vehicles
Person
All
FLOPS (G)
RGB ${}^{\dagger}$
0.790
0.873
0.808
0.559
0.664
0.752
0.891
0.967
0.399
0.000
0.670
45
Thermal ${}^{\dagger}$
0.766
0.835
0.789
0.536
0.654
0.749
0.907
0.971
0.399
0.000
0.661
45
EAEFNet [38]
0.805
0.863
0.813
0.545
0.702
0.802
0.937
0.976
0.481
0.000
0.692
358
CRM [59]
0.816
0.900
0.851
0.616
0.695
0.779
0.923
0.974
0.377
0.206
0.714
310
CMNeXt [77]
0.830 0.900 0.861
0.607
0.718 0.808
0.933
0.980 0.560
0.190
0.740
148
${}^{\dagger}$ DeepLabV3+ (w/ ResNet50 encoder)

### v2

# Caltech Aerial RGB-Thermal Dataset in the Wild

@@ Abstract
We present the first publicly-available RGB-thermal dataset designed for aerial robotics operating in natural environments. Our dataset captures a variety of terrain across the United States, including rivers, lakes, coastlines, deserts, and forests, and consists of synchronized RGB, thermal, global positioning, and inertial data. We provide semantic segmentation annotations for 10 classes commonly encountered in natural settings in order to drive the development of perception algorithms robust to adverse weather and nighttime conditions. Using this dataset, we propose new and challenging benchmarks for thermal and RGB-thermal (RGB-T) semantic segmentation, RGB-T image translation, and motion tracking. We present extensive results using state-of-the-art methods and highlight the challenges posed by temporal and geographical domain shifts in our data. The dataset and accompanying code is available at https://github.com/aerorobotics/caltech-aerial-rgbt-dataset.
@@
**footnotetext: These authors contributed equally to this work.
Figure 1: Left: Our dataset is uniquely designed to improve thermal scene perception for field robots. Right: We provide new benchmarks for thermal-based vision algorithms, including (a) semantic segmentation, (b) image translation, and (c) motion tracking.
@@ 1 Introduction
Field robots rely predominantly on visual cameras, lidar, and radar to perceive their surroundings [16, 18]. These sensors provide robust perception capabilities, but falter in low-light and adverse weather conditions [70]. In contrast, thermal cameras use long-wave infrared wavelengths to capture emitted heat, offering dense, visual information in such conditions where other sensors struggle [19]. Recently, they have been used for multimodal perception in autonomous driving and are increasingly being used to enable nighttime robot autonomy [36, 76, 71, 11, 78, 49, 12, 37]. However, successful integration of thermal cameras requires extensive datasets across diverse settings, preventing widespread adoption in field robotics.
@@
Current thermal datasets focus on urban settings for autonomous driving applications (Fig. 1). They comprise of thermal-only imagery [76, 36] or RGB-Thermal (RGB-T) pairs [17, 22, 71, 41], with some including global positioning (GPS/GNSS) and inertial measurements (IMU) for visual-inertial odometry (VIO) and simultaneous localization and mapping (SLAM) [78, 11, 61]. While comprehensive, they lack data depicting natural settings like rivers and forests, which are typical operating areas for field robotics that perform coastline mapping and bathymetry [4] or activity monitoring during forest fires [30].
@@
Due to the lack of thermal benchmarks for algorithms like semantic segmentation and SLAM, field robots cannot easily operate at night or in adverse conditions, requiring online learning [72, 34] or unsupervised domain adaptation to compensate [20, 32, 71], which still require labeled data for evaluation. As such, enabling thermal perception requires collecting and annotating field-specific data from scratch. However, creating a thermal dataset for field robotics is difficult due to two reasons: First, thermal data is hard to crowdsource and web-scrape due to the nicheness of thermal cameras. Hence, acquiring a diverse dataset for algorithm development requires physically going to distinct field locations, increasing time and cost. Second, operating field robots, especially uninhabited aerial vehicles (UAV), usually requires special permits for each location-of-capture.
@@
In this work, we present the first dataset targeting thermal scene perception for field robotics. The dataset consists of oblique-facing imagery captured from a UAV, supplemented with image sets captured at ground level, and focuses on littoral settings within various desert, forest, and coastal environments across the United States. We contribute new benchmarks for thermal and RGB-T semantic segmentation, RGB-T image translation, and VIO/SLAM algorithms, with unique challenges characterized by temporal and geographical domain shifts for learning-based methods and feature sparsity for motion-tracking algorithms. Aside from the proposed benchmarks, our dataset can also be used towards other applications like visual place recognition and terrain-relative navigation.
@@ 2.1.1 Thermal/RGB-T benchmarks for urban robotics
Current urban thermal and RGB-T datasets primarily focus on autonomous vehicle (AV) technology and surveillance applications. AV-related datasets cover tasks such as object detection [17, 41, 11], semantic segmentation [36, 22, 71, 76], and localization [78, 11, 61]. Some include data from other spectra, benchmarking not only thermal algorithms but also RGB-T [41, 11, 22, 71, 63] and other multispectral models. In contrast, datasets for surveillance applications use fixed cameras [29] and UAVs [67, 63] to detect personnel, vehicles, and other objects. Despite many urban thermal benchmarks, they cannot be directly used to develop algorithms for robots operating in natural environments due to the domain gap.
@@ 2.1.2 Thermal/RGB-T benchmarks for field robotics
Datasets for robots operating in non-urban environments remain limited. MassMIND [49] and PST900 [62] benchmark thermal semantic segmentation for maritime and subterranean robotic environments, respectively. WIT-UAS [30] provides UAV-borne thermal images for object detection in wildfire-prone environments. One work [34] proposes an aerial RGB-T dataset for thermal water segmentation in littoral areas but only provides 1272 labels consisting of a single water class. Benchmarks like [3, 7] depict natural environments but are targeted towards wildlife tracking. To address this gap, we present an aerial RGB-T dataset which can be used to benchmark multiple tasks, including semantic segmentation, RGB-T image translation, and localization, across diverse natural environment robotic applications. Specifically, we build upon the dataset from [34] by extending the number of semantic classes from 1 to 10 and adding additional RGB-T data captured from air and ground, resulting in a total of 4195 semantic segmentation annotations. In comparison to existing thermal/RGB-T datasets, our dataset can enable nighttime field robotic applications like remote sensing research, geological sampling, precision agriculture, and wilderness search and rescue.
@@ 2.1.3 UAV-based semantic segmentation datasets
Aside from [34], existing UAV-captured semantic segmentation datasets [48, 10, 45, 6, 47, 63] cover urban environments (Tab. 1). With the exception of [47] which provides 400 thermal image samples, all of these datasets contain only RGB data for semantic segmentation. While they provide images with relevant view angles, the modality gap and focus on urban content makes them impractical for aerial field robotic applications.
Table 1: Comparison of datasets captured from aerial platforms. Camera angle of 90°is nadir-pointing. See Tab. S1 in supplement for comparison against thermal datasets captured from surface and ground vehicles.
Dataset Task RGB Thermal GPS IMU Setting Location # Samples Camera pose
VisDrone [83] Obj. Det. Urban China 10,209 Variable
HIT-UAV [67], NII-CU [63] Obj. Det. Urban China, Japan 2,898 - 5,880 30°-90°, 60-130 m
WIT-UAS [30] Obj. Det. Forest USA 6,951 30°, 90°, < 120 m
BIRDSAI [3] Obj. Track. Savannas S. Africa 62,000 Off-nadir, 60-120 m
AeroScapes [48], UAVid [45], [10][6], Swiss/Okutama [63] Sem. Seg. Urban China, Japan, Switzerland 191 - 3269 30 - 90°, < 100 m
Semantic Drone Dataset [47] Sem. Seg. Urban Germany 400 90°, < 30m
Ours Sem. Seg. River, Coast, Lake, Mountain, Desert USA 4,195 20°, 45°, < 120 m
@@ 2.2.4 Thermal VIO and SLAM
3 The Caltech Aerial RGB-Thermal Dataset
@@ 3.2.3 Semantic segmentation annotation
To promote labeling consistency, annotators preread descriptions of each trajectory, which contained a list of classes present, labeled examples provided by the authors, and GPS coordinates (for Google Earth look-up). Annotators were then given an RGB-T image pair with trajectory metadata and asked to label the thermal frame. Annotations were reviewed by authors, with inconsistent or incorrect labels sent back. 3 rounds of annotation were done, with the final and initial set of labels differing by 0.08 mIoU. Classes most frequently re-annotated were bare ground, rocky terrain, road, trees, and vehicles.
@@ 3.3.3 Geographical split
    Region-based: This is based on the area of the United States where the data was captured. Distinct regions include California (CA), Kentucky (KY), and North Carolina (NC), with train/val/test created in the same way as above.
@@ 4 Experiments
Model Bare ground Rocky terrain Develop. struct. Road Shrubs Trees Sky Water Vehicles Person Stuff Object All Orin (GPU) NUC (CPU) FLOPS (G)
FastSCNN [53] 0.807 0.907 0.728 0.631 0.694 0.776 0.956 0.965 0.384 0.058 0.808 0.221 0.690 124.2 85.6 1.0
MobileNetV3-S 0.75 [25] 0.802 0.914 0.676 0.633 0.704 0.814 0.955 0.976 0.383 0.111 0.809 0.247 0.697 75.2 24.0 3.3
MobileNetV3-L 0.75 0.801 0.914 0.701 0.615 0.713 0.793 0.959 0.976 0.432 0.148 0.809 0.290 0.705 59.0 15.5 4.8
MobileViTV2 0.50 [46] 0.822 0.913 0.759 0.650 0.694 0.781 0.954 0.970 0.222 0.030 0.818 0.126 0.680 60.1 10.7 5.5
EfficientViT-B0 [5] 0.825 0.917 0.777 0.618 0.710 0.793 0.961 0.965 0.495 0.191 0.821 0.343 0.725 51.2 5.5 3.9
EfficientNet-Lite0 [68] 0.819 0.908 0.769 0.601 0.709 0.805 0.962 0.968 0.380 0.079 0.818 0.229 0.700 47.7 12.8 7.2
EfficientNet-Lite2 0.825 0.920 0.765 0.614 0.698 0.805 0.958 0.971 0.386 0.118 0.820 0.252 0.706 38.9 11.2 9.4
Segformer-B0 [74] 0.804 0.910 0.690 0.624 0.696 0.788 0.960 0.970 0.195 0.000 0.805 0.097 0.664 38.3 5.9 10.2
Segformer-B1 0.814 0.909 0.773 0.603 0.713 0.799 0.960 0.965 0.366 0.000 0.817 0.183 0.690 30.3 3.5 19.9
ResNet18 [23] 0.810 0.905 0.766 0.618 0.711 0.807 0.959 0.966 0.431 0.158 0.818 0.294 0.713 69.3 12.4 22.4
ResNet50 0.819 0.916 0.747 0.658 0.711 0.799 0.961 0.974 0.361 0.182 0.823 0.272 0.713 29.9 6.1 45.4
ResNeXt50 [75] 0.825 0.919 0.773 0.653 0.709 0.794 0.964 0.976 0.436 0.137 0.827 0.287 0.719 23.0 5.6 45.5
ConvNext-T [44] 0.799 0.909 0.719 0.608 0.706 0.808 0.961 0.970 0.363 0.003 0.810 0.183 0.685 25.8 3.3 47.4
ConvNext-S 0.810 0.913 0.771 0.603 0.718 0.810 0.965 0.964 0.492 0.048 0.819 0.270 0.709 19.1 2.3 75.0
ConvNext-B 0.810 0.921 0.697 0.611 0.706 0.803 0.963 0.972 0.348 0.112 0.811 0.230 0.694 14.6 1.4 130.5
ConvNext-B (CLIP) [55] 0.813 0.918 0.683 0.632 0.718 0.813 0.965 0.972 0.517 0.137 0.814 0.327 0.717 14.6 1.4 130.4
ConvNext-B (CLIP) ❄ 0.773 0.887 0.713 0.487 0.656 0.756 0.946 0.951 0.158 0.019 0.771 0.089 0.635 14.6 1.4 130.4
DINOv2 [50] (linear head) 0.800 0.907 0.681 0.606 0.693 0.796 0.956 0.967 0.375 0.079 0.801 0.227 0.686 15.2 2.1
DINOv2 (linear head) ❄ 0.705 0.844 0.619 0.419 0.558 0.725 0.922 0.920 0.356 0.158 0.714 0.257 0.623 15.2 2.1
DINOv2 (nonlin. head) 0.810 0.916 0.708 0.635 0.700 0.819 0.959 0.973 0.399 0.119 0.815 0.259 0.704 15.2 2.1
DINOv2 (nonlin. head) ❄ 0.796 0.903 0.732 0.606 0.691 0.794 0.954 0.963 0.471 0.149 0.805 0.310 0.706 15.2 2.1
FTNet [51] 0.755 0.867 0.635 0.576 0.643 0.653 0.787 0.947 0.234 0.024 0.733 0.129 0.613 11.1 1.4 100.5
Thermal-specific model  ❄ Frozen encoder
@@ 4.1.3 Geographically-partitioned benchmarks
(Tab. 2) on geographically-partitioned sets to evaluate intra-thermal generalization. Given the cost and effort of collecting diverse aerial thermal data of field settings, we aim to determine the minimum amount of annotated data required to effectively deploy. Using EfficientViT-B0, we performed two experiments to study generalization across difference geographic regions and terrain types.
@@
River Lake Coast
River 0.668 0.196 0.117 0.327
Lake 0.196 0.364 0.061 0.207
Coast 0.127 0.119 0.522 0.256
All 0.646 0.415 0.493 0.518
All + ground 0.658 0.416 0.520 0.532
@@ 4.1.4 Temporally-partitioned benchmark
Model Bare ground Rocky terrain Devel. struct. Road Shrubs Trees Sky Water Vehicles Person All FLOPS (G)
RGB Only 0.790 0.873 0.808 0.559 0.664 0.752 0.891 0.967 0.399 0.000 0.670 45
Thermal Only 0.766 0.835 0.789 0.536 0.654 0.749 0.907 0.971 0.399 0.000 0.661 45
EAEFNet [39] 0.805 0.863 0.813 0.545 0.702 0.802 0.937 0.976 0.481 0.000 0.692 358
CRM [60] 0.816 0.900 0.851 0.616 0.695 0.779 0.923 0.974 0.377 0.206 0.714 310
CMNeXt [79] 0.830 0.900 0.861 0.607 0.718 0.808 0.933 0.980 0.560 0.190 0.740 148
DeepLabV3+ (w/ ResNet50 encoder)